OpenStudy (anonymous):

Divide the following polynomial (6x^3+11x^2-4x-4)/(3x-2)

OpenStudy (anonymous):

well first simplify (6x^3 + 11x^2 - 4x - 4) / (3x - 2) (6x^3 + 11x^2) - (4x - 4) / (3x - 2) x^2(6x + 11) - 4 (x - 1) / (3x - 2) I don't think you can divide the polynomial it does not work out.

OpenStudy (anonymous):

@friedRice but its long division

OpenStudy (anonymous):

yeah I changed it into long division and it did not work

OpenStudy (anonymous):

(here is what they say) You have to use a method called polynomial long division. I found this question divides with no remainder, My answer was 2x^2+5x+2.

OpenStudy (anonymous):

(6x + 11)(x + 1)(x^2 - 4) / (3x -2)

OpenStudy (anonymous):

that's long division

OpenStudy (anonymous):

check

OpenStudy (anonymous):

(x 2 - 4)/ (x - 1) that's not long division but my answer doesn't match to put on grid

OpenStudy (anonymous):

different problem

OpenStudy (anonymous):

take look that's how it should looks like

OpenStudy (anonymous):

(x^2 -4) / (x - 1) (x + 2)(x - 2) / (x - 1) well I don't know but if yahoo does it like that you will get something

OpenStudy (anonymous):

and the other people did not divide right

OpenStudy (anonymous):

look 3x - 2 | 6x^3+11x^2-4x-4 | 2x^2 + 5x +2 6x^3 - 4x^2 (-)

OpenStudy (anonymous):

carefully

OpenStudy (anonymous):

did you found the answer?

OpenStudy (anonymous):



OpenStudy (anonymous):

from dividing 3x - 2

OpenStudy (anonymous):

@FriedRice yeah so what you think witch answer?

OpenStudy (anonymous):

@FriedRice well 11/5 = 5.5 they just removed point i think

OpenStudy (anonymous):

lets ask a mentor

OpenStudy (anonymous):

who is mentor?

OpenStudy (anonymous):

lets check

OpenStudy (anonymous):

be right back

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@amistre64 @Hero @ash2326

OpenStudy (amistre64):

well, long division is soo passe; synthetic division is a much better process

OpenStudy (anonymous):

Divide the following polynomial (6x^3+11x^2-4x-4)/(3x-2)

OpenStudy (amistre64):

6x^3 + 11x^2 - 4x - 4 0 4 10 4 2/3 6 15 6 0

OpenStudy (amistre64):

\[\frac{6x^3+11x^2-4x-4}{3x-2}\] \[\frac13\frac{6x^3+11x^2-4x-4}{x-\cfrac23}\]

OpenStudy (amistre64):

2x^2 + 5x + 2

OpenStudy (anonymous):

im confused how you got 5

OpenStudy (amistre64):

15/3 = 5

OpenStudy (anonymous):

where you got 15?

OpenStudy (anonymous):

nvmyou added 4

OpenStudy (amistre64):

i used sythetic division to make the process simpler

OpenStudy (anonymous):

amister64 do you know any online calculator that solves those kind problems?

OpenStudy (amistre64):

i just use wolfram if i need to dbl chk stuff

OpenStudy (anonymous):

i did it to but it seems showing me diffrent answers

OpenStudy (anonymous):

(x^2 - 4)/(x - 1) i tryed put this on wolfram and can't find the answer

OpenStudy (anonymous):

didn't find the answer to match this look on screen

OpenStudy (amistre64):

if we do long hand (6x^2 + 15x + 6 ) /3 ----------------------- x-2/3 | 6x^3 + 11x^2 - 4x - 4 -(6x^3 -4x^2) -------------- 15x^2 - 4x -(15x^2-10x) ------------- 6x - 4 -(6x - 4) --------- 0 (6x^2 + 15x + 6 ) /3 = 2x^2 +5x + 2

OpenStudy (anonymous):

(x^2 - 4)/(x - 1) i tryed put this on wolfram and can't find the answer

OpenStudy (anonymous):

help me with this one

OpenStudy (amistre64):

x^2 + 0x - 4 0 1 1 1 1 1 -3 \[{x+1-\frac{3}{x-1}}{}\]

OpenStudy (anonymous):

can you show how did you put on wolfram

OpenStudy (anonymous):

i had same thing but where you see the answer?

OpenStudy (amistre64):

if you type simplify in front of it you get alternate forms

OpenStudy (amistre64):

there is no "one" alternative

OpenStudy (anonymous):

the answer you gave me i can't find on wolfram umm can you screen and circle with red the answer

OpenStudy (amistre64):

its the first one ... remember that addition of terms is not specific to position 1 + 2 + 3 is equal to the same setup as: 3 + 1 + 2

OpenStudy (anonymous):

look

OpenStudy (anonymous):

so you moved 1 from the end?

OpenStudy (amistre64):

i tend to put the remainder at the end of a quotient string ... the wolf seems to feel it more appropriate to place a constant at the end

OpenStudy (anonymous):

where is the answer for this one

OpenStudy (amistre64):

all three results that they show are equally valid. You are pretty much asking if: \(1.75\), or \(1\cfrac34\), or \(\left(1+\cfrac34\right)\), or \(\left(\cfrac34+1\right)\), or \(\cfrac{7}{4}\) ... is more correct

OpenStudy (anonymous):

(a^ (2n) - a^n - 6)/(a^n + 8) <---- for this problem?

OpenStudy (amistre64):

yes, and for the rest of your problems as well

OpenStudy (amistre64):

is can be written in several different ways .... which way is the "correct" way? that depends on whoever wants to grade the work

OpenStudy (amistre64):

it gives you 4 different ways to write it out ... they are all equal to each other

OpenStudy (anonymous):

here how it should look in my problem i don't know witch one to choice

OpenStudy (amistre64):

the first one then

OpenStudy (anonymous):

when i start putting symbols in a^2+[ ] [ ] their is 2 spaces before other parts begin

OpenStudy (amistre64):

then since you cant possible fit -9 at the end of it ... plug it in the middle of it

OpenStudy (anonymous):

but a^2+-9 you think that will work?

OpenStudy (amistre64):

of course not .... a^2 - 9 + ________

OpenStudy (anonymous):

oh ok got it:)so confusing hate polynomials now

OpenStudy (amistre64):

its not really the polys fault, its that rediculous program you have to fight with ....

OpenStudy (anonymous):

wolf?

OpenStudy (amistre64):

yeah, the wolf is making you fill in boxes in a specific manner ....

OpenStudy (anonymous):

now all fits :) so i need to rember if at the end has numbers such as -9 i put in middle right?

OpenStudy (amistre64):

i would, it looks like your program there does what i do, tag the remainder on the end.

OpenStudy (anonymous):

alright thank you for helping me out gave you medal :)

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

have nice day